# Causal machine learning for heterogeneous treatment effects in the presence of missing outcome data

### v1

@@ Abstract
When estimating heterogeneous treatment effects, missing outcome data can complicate treatment effect estimation, causing certain subgroups of the population to be poorly represented. In this work, we discuss this commonly overlooked problem and consider the impact that missing at random (MAR) outcome data has on causal machine learning estimators for the conditional average treatment effect (CATE). We then propose two de-biased machine learning estimators for the CATE, the mDR-learner and mEP-learner, which address the issue of under-representation by integrating inverse probability of censoring weights into the DR-learner and EP-learner respectively. We show that under reasonable conditions, these estimators are oracle efficient, and illustrate their favorable performance through simulated data settings, comparing them to existing CATE estimators, including comparison to estimators which use common missing data techniques. Guidance on the implementation of these estimators is provided and we present an example of their application using the ACTG175 trial, exploring treatment effect heterogeneity when comparing Zidovudine mono-therapy against alternative antiretroviral therapies among HIV-1-infected individuals.
@@ 1 Introduction
In recent years, many estimators for the CATE have been proposed, with the focus turning towards non-parametric estimators, using machine learning (ML) to estimate complex functions of high dimensional data (Künzel et al., 2019; Kennedy, 2023; Nie and Wager, 2021; van der Laan et al., 2024). Of these estimators, many require that the training data be fully-observed and no data be missing. In this paper, we explore how causal ML estimators can be constructed for the CATE when outcome data is missing at random (MAR), i.e., outcome missingness depends only on the treatment and observed baseline covariates.
MAR outcome data occurs frequently in practice, typically arising when individuals are lost to follow up. When it occurs, the observed data may no longer represent the target population, and subgroups that have high levels of drop-out can be under-represented. This presents a challenge for existing non-parametric CATE estimators, which do not address this under-representation and are prone to producing biased estimates of the CATE within these under-represented subgroups. To overcome this, some authors propose using CATE estimators in combination with established missing data techniques, such as imputing missing outcomes (Groenwold et al., 2014; Berrevoets et al., 2023), or re-weighting the population using inverse probability of censoring weights (IPCW) (Robins et al., 1994; Gonzalez Ginestet et al., 2021). However, when implementing these approaches using non-parametric (machine learning) techniques, the inherent slow convergence of ML algorithms (Hastie, 2009) can introduce errors into the IPCW or imputation predictions, which subsequently propagate through to the CATE estimates.
We explore this issue further, discussing the shortfalls of the existing CATE estimators when handling missing outcome data (Section 2). We then demonstrate how IPCWs can be robustly incorporated into CATE estimators, proposing two non-parametric CATE estimators, the mDR-learner and mEP-learner, which are extensions of the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024) respectively (Section 3). We demonstrate the empirical performance of our proposed estimators through a simulation study (Section 4) and illustrate their application to data from the ACTG175 trial (Hammer et al., 1996), where we explore treatment effect heterogeneity across baseline covariates, comparing Zidovudine mono-therapy against alternative antiretroviral therapies among HIV-1-infected subjects (Section 5). We then discuss the potential for further extensions in Section 6.
@@ 2.1 Setting
We define a sample of $n$ independent and identically distributed (i.i.d.) observations drawn from $O=(Z,A,CY)$, where $A$ is a binary treatment, $A\in\{0,1\}$ and $Y$ is a continuous outcome. In this setting, we allow outcome data to be missing and define $C$ as an indicator of outcome missingness, $C\in\{0,1\}$, with $C=1$ indicating that an individual’s outcome is non-missing. We then define $Z$ to be a set of baseline covariates which contains all confounders between the treatment $A$ and the outcome $Y$, and also contains sufficient information for missing outcomes to be assumed MAR (i.e. after controlling for $Z$ and $A$, the missingness indicator $C$ and outcome $Y$ are independent).
The focus is on learning the CATE, $\theta(x)=\operatorname{\mathbb{E}}[Y(1)|X=x]-\operatorname{\mathbb{E}}[Y(0)|X=x]$, which is conditional on $X\subseteq Z$, where $X$ are the covariates for which heterogeneity is of interest.
@@ 2.2 Estimating the CATE using ML - No missing outcome data
Let us first consider how causal ML estimators are constructed when no outcome data is missing. Under the standard causal assumptions of (A1) consistency, (A2) no unmeasured confounding, $Y(a)\perp\!\!\!\perp A|Z$ for $a\in\{0,1\}$ and (A3) positivity of treatment exposure, $10<P(A=1|Z)<$ 1 (with probability 1) (Pearl et al., 2016), the CATE can be identified as:
@@
ML can be used to directly estimate the conditional expectations or probabilities (nuisance functions) found in eq. (1) or eq. (2), with these estimators commonly referred to as plug-in estimators. One common example of a plug-in estimator is the T-learner (Künzel et al., 2019), which requires $X=Z$ and estimates the conditional expectations from eq. (1) in the subsets of individuals who are treated/untreated, $\mu^{1}(Z)=\operatorname{\mathbb{E}}[Y|A=1,Z]$ and $\mu^{0}(Z)=\operatorname{\mathbb{E}}[Y|A=0,Z]$, taking their difference to obtain CATE estimates. Alternatively, plug-in estimators based on eq. (2) are often referred to as inverse probability of treatment weight (IPTW) estimators, as they require the estimation of the propensity score, $\pi(Z)=P[A=1|Z]$ (Kennedy, 2023).
@@
While such ML-based plug-in estimators are simple to understand and implement, they are also prone to bias. This arises as ML algorithms use regularization to avoid over-fitting to the training data, reducing the rate at which these estimators converge towards the true parameter. This introduces non-negligible bias when fit using finite samples, with the errors in these nuisance function estimates (e.g. $\hat{\mu}^{1}(Z),\hat{\mu}^{0}(Z)$ or $\hat{\pi}(Z)$) directly propagating through to the estimates of the CATE, known as “plug-in bias” (Morzywolek et al., 2023). Furthermore, plug-in estimators which estimate the conditional expectations of the outcome, such as the T-learner, do not ensure an optimal bias-variance trade-off is made for the CATE, as they optimize predictions for their outcome functions rather than for the CATE itself (Kennedy, 2023). See Appendix A for an illustrative example.
@@
Estimators that target the CATE directly and that are less sensitive to nuisance model estimation avoid the issues noted above. Estimators of this form are typically constructed using the efficient influence function (EIF) of the estimand of interest, where the EIF represents how sensitive a measure of prediction error is to changes in the data generating distribution (Chernozhukov et al., 2017; Ichimura and Newey, 2022). The EIF offers a useful tool for constructing estimators, as it allows estimators to be designed which are approximately insensitive to small changes in its nuisance functions, meaning estimation errors in the outcome functions, $\mu^{1}(Z)$ and $\mu^{0}(Z)$, or propensity score, $\pi(Z)$, affect errors in the estimates of the target function only through their product. Unfortunately the EIF of the CATE is generally not well defined (i.e., has infinite variance) whenever it depends on continuous variables (Hines et al., 2022). Instead, estimators which achieve these properties can be constructed based on the EIF of a well-chosen loss function, in this case, a measure of counterfactual prediction error (Morzywolek et al., 2023). Two estimators constructed in this way are the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024).
@@
The DR-learner is a model-agnostic method for estimating the CATE, meaning the user can choose any estimation strategy, including data adaptive methods, when estimating any functions within it. It is derived using a two-step procedure (Kennedy, 2023), with the first step calculating pseudo-outcomes, $Y_{DR}$, using the EIF of the mean square error (MSE) for the CATE:
@@
where $\hat{\mu}^{A}(Z)=A\cdot\hat{\mu}^{1}(Z)+(1-A)\cdot\hat{\mu}^{0}(Z)$. The second step then learns the CATE by regressing the pseudo-outcomes on the covariate set in which heterogeneity is of interest, $X$ (See Appendix B for the full algorithm).
The DR-learner is an example of a one-step IF based estimator and is derived by considering the components of the mean squared error (MSE) of the CATE that depend on $\theta(X)$,
$\psi_{CATE}=\operatorname{\mathbb{E}}\left[\theta(X)^{2}-2\theta(X)(\mu^{1}(Z)-\mu^{0}(Z))\right].$ (4)
It then uses the EIF of this risk function,
@@
and defines its pseudo-outcomes such that the sample average of the drift term (second term in eq. (5)) goes to 0. These pseudo-outcomes closely resemble the EIF from the average treatment effect (ATE) and remove the plug-in bias from the plug-in estimator, $\hat{\mu}^{1}(Z)-\hat{\mu}^{0}(Z)$, by adding the drift term, a weighted error, to these estimates. By defining the pseudo-outcomes in this way, the DR-learner ensures that the gradient of the MSE risk function, with respect to the CATE, is less sensitive to errors in its nuisance functions. It also allows the DR-learner to achieve oracle efficiency, meaning that when the outcome models and propensity score model converge to the truth at sufficiently fast rates, it performs asymptotically as if the nuisance functions were known (Kennedy, 2023). This allows data adaptive techniques to be used to estimate the nuisance functions, provided they converge sufficiently fast to the truth, without fear of errors propagating through to the CATE. However, we note that the convergence rate of the DR-learner will still depend on the convergence rate of its second-stage pseudo-outcome regression, which will vary itself depending on the complexity of the CATE.
Despite its desirable properties, the DR-learner’s pseudo-outcomes are sensitive to extreme propensity scores. This occurs as the DR-learner uses IPTWs when calculating its pseudo-outcomes (first term of (3)), meaning that when estimates of the propensity score are near 0/1, the pseudo-outcomes can grow infinitely large. To prevent this, an alternative approach known as infinite-dimensional targeting (iTMLE) can be used (Luedtke et al., 2017; Vansteelandt and Morzywołek, 2023; van der Laan et al., 2024). This technique has been developed for counterfactual outcome prediction (Vansteelandt and Morzywołek, 2023), and tends to have less sensitivity to extreme propensity scores as its targeted learning framework can moderate the impact of outlier propensity score estimates. In this paper we discuss a very similar approach known as the EP-learner (efficient plug-in learner) (van der Laan et al., 2024), which uses iTMLE in the context of causal contrasts, including CATE estimation.
@@
The targeting step then works by regressing the outcomes $Y$ on the sieve basis, $\varphi(X)$, in a weighted linear regression, with an offset $\mu^{A}(Z)$, and weight, $\hat{H}(A,Z)=\frac{A}{\hat{\pi}(Z)}+\frac{1-A}{\left(1-\hat{\pi}(Z)\right)}$, suggested by the EIF of the risk function, eq. (5). After fitting this model, efficient plug-in estimates, $\hat{\mu}^{0*}(Z)$ and $\hat{\mu}^{1*}(Z)$ are obtained by adding/subtracting the estimated linear predictor from this model to the plug-in outcome estimates (See Appendix C for the full algorithm). As the updated outcome predictions are defined such that the sample average of the drift term in eq. (5) converges to 0, the EP-learner achieves the same oracle efficiency properties as the DR-learner. Additionally, by using an iTMLE approach, extreme pseudo-outcomes are less common and the CATE estimates obtained by the EP-learner are expected to be more stable (van der Laan et al., 2024).
In addition, both the EP-learner and DR-learner require that the data which used within the second stage optimization is i.i.d. To achieve this, sample splitting is used when estimating the nuisance functions, preventing correlations from being introduced into the pseudo-outcomes. In this paper, we implement sample splitting by using a K-fold cross-validation process which works by splitting the data into K-folds, training each nuisance model on all folds except one, then obtaining predictions for the individuals in the remaining fold. This is repeated until nuisance function estimates are obtained for all individuals.
@@ 2.3 Estimating the CATE using ML - With missing outcome data
Now let us consider how missing outcome data impacts the existing estimators. To identify the CATE when outcome data is MAR, we require assumptions A1-A3 from Section 2.2 and two additional assumptions on the missingness mechanism: (A4) Outcomes are MAR, $Y\perp\!\!\!\perp C~{}|~{}A,Z$; (A5) Positivity of outcomes being non-missing, $0<P(C=1|A,Z)$, with probability 1.
Under assumptions A1-A5, the CATE can then be identified as eq. (6) and eq. (7).
@@
The no unmeasured confounding assumption (A2) and MAR assumption (A4) need not be conditional on the same covariates, however, for simplicity, we define both assumptions to be conditional on the same set of covariates, $Z$.
@@
Using equations (6) and (7) we see that when all of the covariates in $Z$ are present, the CATE can be estimated using the observed data. This means the existing CATE estimators, such as the T-learner, DR-learner and EP-learner can produce asymptotically unbiased estimates of the CATE by restricting their analyses to complete cases, and adjusting for $Z$ through their outcome models $\mu^{0}(Z)=\operatorname{\mathbb{E}}[Y|A=0,C=1,Z]$ and $\mu^{1}(Z)=\operatorname{\mathbb{E}}[Y|A=1,C=1,Z]$. We refer to this approach as an “available case analysis”, and note that while the outcome regressions are limited to complete cases, the propensity score models should still be estimated using the full dataset.
@@
Available case analyses offer the simplest way of estimating the CATE in the presence of MAR outcome data, however, asymptotic unbiasedness does not mean these methods perform well when fit using finite samples, and available case analyses can be inefficient for a range of reasons. Firstly, as the outcome regressions are restricted to those with complete cases, information is ignored for individuals who have a missing outcome. Equally, when the outcome models are fit using ML, they will be prone to over-smoothing in the subsets of the population which experience high levels of outcome missingness. This can lead available case analyses to miss complex non-linear CATEs, over-smoothing outcome predictions due to the missing data, or can even lead to estimators identifying heterogeneity where non-exists, for instance when the missingness only occurs within one treatment arm. Specific examples of data generating processes (DGPs) where this can occur are presented in Section 4.
@@
Because of the limitations of available case analyses, some authors choose to address missing outcome data by utilizing missing data techniques. One common missing data technique is to impute outcomes, replacing the missing outcomes with outcome predictions gained from an imputation model, $\hat{\operatorname{\mathbb{E}}}[Y|C=1,A,Z]$ (Groenwold et al., 2014; Berrevoets et al., 2023). These are easy to implement when using any of the existing non-parametric CATE estimators, as the estimators can be run using the imputed, complete dataset. However, as outcome imputations require the estimation of an additional nuisance model, plug-in bias can be introduced to the existing estimators, with errors in the outcome imputations propagating through to the estimates of the CATE (even for estimators which have oracle efficiency in the complete data setting). Alternatively, other authors suggest addressing missing outcome data by re-weighting observed individuals based on their probability of having a non-missing outcome, $G(A,Z)=P[C=1|A,Z]$ (Robins et al., 1994; Gonzalez Ginestet et al., 2021). These weights can be used to estimate the CATE using a similar approach to the one Kennedy (2023) uses to incorporate IPTWs, weighting observed individuals to create pseudo-outcomes, then regressing the pseudo-outcomes against $X$ to estimate the CATE. However, these estimators will also suffer from plug-in bias, as errors in the IPCWs will propagate through to the CATE estimates. For this reason, we explore how to incorporate IPCWs in both the DR-learner and EP learner such that the resulting learners are robust to estimation of the missingness model $G(A,Z)$.

### v2

@@ Abstract
When estimating heterogeneous treatment effects, missing outcome data can complicate treatment effect estimation, causing certain subgroups of the population to be poorly represented. In this work, we discuss this commonly overlooked problem and consider the impact that missing at random (MAR) outcome data has on causal machine learning estimators for the conditional average treatment effect (CATE). We propose two de-biased machine learning estimators for the CATE, the mDR-learner and mEP-learner, which address the issue of under-representation by integrating inverse probability of censoring weights into the DR-learner and EP-learner respectively. We show that under reasonable conditions, these estimators are oracle efficient, and illustrate their favorable performance through simulated data settings, comparing them to existing CATE estimators, including comparison to estimators which use common missing data techniques. We present an example of their application using the GBSG2 trial, exploring treatment effect heterogeneity when comparing hormonal therapies to non-hormonal therapies among breast cancer patients post surgery, and offer guidance on the decisions a practitioner must make when implementing these estimators.
@@ 1 Introduction
In recent years, many estimators for the CATE have been proposed, with the focus turning towards non-parametric estimators, using machine learning (ML) to estimate complex functions of high dimensional data (Künzel et al., 2019; Kennedy, 2023; Nie and Wager, 2021; van der Laan et al., 2024). Of these estimators, each require that the training data be fully-observed and no data be missing. In this paper, we relax this requirement and propose two novel CATE estimators, the mDR-learner and mEP-learner, which demonstrate how causal ML estimators can be constructed when outcome data is missing at random (MAR).
MAR outcome data occurs frequently in practice, typically arising when individuals are lost to follow up. When it occurs, the observed data may no longer represent the target population, and subgroups that have high levels of drop-out can be under-represented. This presents a challenge for existing non-parametric CATE estimators, which do not address this under-representation and are prone to producing biased estimates of the CATE within these under-represented subgroups. To overcome this, some authors propose using CATE estimators in combination with established missing data techniques, such as imputing missing outcomes (Groenwold et al., 2014; Berrevoets et al., 2023), or re-weighting the population using inverse probability of censoring weights (IPCW) (Robins et al., 1994; Gonzalez Ginestet et al., 2021). However, when implementing these approaches using non-parametric, machine learning techniques, the inherent slow convergence of ML algorithms can introduce errors into the IPCW/imputation predictions, which then propagate through to the CATE estimates.
Our work aims to overcome this issue, with the mDR-learner and mEP-learner robustly incorporating IPCWs into the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024) respectively. We show that these estimators are oracle efficient under reasonable conditions and demonstrate their empirical performance through a simulation study (Section 4). We then illustrate their application, exploring treatment effect heterogeneity within the GBSG2 trial (Section 5) and finish by discussing potential extensions (Section 6).
@@ 2.1 Setting
We define a sample of $n$ independent and identically distributed (i.i.d.) observations drawn from $O=(Z,A,CY)$, where $A$ is a binary treatment, $A\in\{0,1\}$ and $Y$ is a continuous outcome. In this setting, we allow outcome data to be missing and define $C$ as an indicator of outcome missingness, $C\in\{0,1\}$, with $C=1$ indicating that an individual’s outcome is non-missing. We then define $Z$ to be a set of baseline covariates which contains all confounders between the treatment $A$ and the outcome $Y$, and also contains sufficient information for missing outcomes to be assumed MAR (i.e. after controlling for $Z$ and $A$, the missingness indicator $C$ and outcome $Y$ are independent). Finally, we focus on learning the CATE, $\theta(x)=\operatorname{\mathbb{E}}[Y(1)|X=x]-\operatorname{\mathbb{E}}[Y(0)|X=x]$, which is conditional on $X\subseteq Z$, where $X$ are the covariates for which heterogeneity is of interest.
@@ 2.2 Estimating the CATE using ML - No missing outcome data
We first consider how causal ML estimators are constructed when no outcome data is missing. Under the standard causal assumptions of (A1) consistency, (A2) no unmeasured confounding, $Y(a)\perp\!\!\!\perp A|Z$ for $a\in\{0,1\}$ and (A3) positivity of treatment exposure, $10<P(A=1|Z)<$ 1 (with probability 1) (Pearl et al., 2016), the CATE can be identified as:
@@
ML can be used to estimate the conditional expectations/probabilities (nuisance functions) found in eq. (1) or eq. (2), with these estimators commonly referred to as plug-in estimators. One common example of a plug-in estimator is the T-learner (Künzel et al., 2019), which requires $X=Z$ and estimates the conditional expectations from eq. (1) in the subsets of individuals who are treated/untreated, $\mu^{1}(Z)=\operatorname{\mathbb{E}}[Y|A=1,Z]$ and $\mu^{0}(Z)=\operatorname{\mathbb{E}}[Y|A=0,Z]$, taking their difference to obtain CATE estimates. Alternatively, plug-in estimators based on eq. (2) are often referred to as inverse probability of treatment weight (IPTW) estimators, as they require the estimation of the propensity score, $\pi(Z)=P[A=1|Z]$ (Kennedy, 2023).
@@
While such ML-based plug-in estimators are simple to understand and implement, they are also prone to bias. This arises as ML algorithms use regularization to avoid over-fitting to the training data, reducing the rate at which these estimators converge towards the true parameter. This introduces non-negligible bias when fit using finite samples, with the errors in these nuisance function estimates (e.g. $\hat{\mu}^{1}(Z),\hat{\mu}^{0}(Z)$ or $\hat{\pi}(Z)$) directly propagating through to the estimates of the CATE, known as “plug-in bias” (Morzywolek et al., 2023). Furthermore, the T-learner does not ensure an optimal bias-variance trade-off is made for the CATE, as it optimizes predictions for their outcome functions rather than for the CATE itself (Kennedy, 2023). See Appendix A for an illustrative example.
@@
Estimators that target the CATE directly and that are less sensitive to nuisance model estimation avoid these issues. Estimators of this form are typically constructed using the efficient influence function (EIF) of the estimand of interest, where the EIF represents how sensitive a measure of prediction error is to changes in the data generating distribution. The EIF offers a useful tool for constructing estimators, allowing estimators to be designed which are approximately insensitive to small changes in its nuisance functions, meaning estimation errors in the outcome functions or propensity score affect errors in the estimates of the target function only through their product. Unfortunately, the EIF of the CATE is generally not well defined (i.e., has infinite variance) whenever it depends on continuous variables (Hines et al., 2022). Instead, estimators which achieve these properties can be constructed based on the EIF of a well-chosen loss function, in this case, a measure of counterfactual prediction error (Foster and Syrgkanis, 2023; Morzywolek et al., 2023). Two estimators constructed in this way are the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024).
@@
The DR-learner is a model-agnostic CATE estimator, meaning the user can choose any estimation strategy, including data adaptive methods, when estimating any functions within it. It is derived using a two-step procedure (Kennedy, 2023), with the first step calculating pseudo-outcomes, $Y_{DR}$, using the EIF of the mean square error (MSE) for the CATE:
@@
where $\hat{\mu}^{A}(Z)=A\cdot\hat{\mu}^{1}(Z)+(1-A)\cdot\hat{\mu}^{0}(Z)$. The second step then learns the CATE by regressing the pseudo-outcomes on the covariate set in which heterogeneity is of interest, $X$. See Appendix B for the full algorithm. By using a pseudo-outcome regression, the DR-learner targets the CATE directly, and can benefit from faster convergence rates when the CATE is smooth. Additionally, these pseudo-outcomes are derived by considering the components of the mean squared error (MSE) of the CATE that depend on $\theta(X)$, $\psi_{CATE}=\operatorname{\mathbb{E}}\left[\theta(X)^{2}-2\theta(X)(\mu^{1}(Z)-\mu^{0}(Z))\right]$, using the EIF of this risk function,
@@
to define its pseudo-outcomes such that the sample average of the drift term (second term in eq. (4)) goes to 0. These pseudo-outcomes closely resemble the EIF from the average treatment effect (ATE) and consist of the plug-in estimates and a weighted error term. By defining the pseudo-outcomes in this way, the DR-learner ensures that the gradient of the MSE risk function, with respect to the CATE, is less sensitive to errors in its nuisance functions. It also allows the DR-learner to achieve oracle efficiency, meaning that when the product of the convergence rates for the outcome predictions and propensity score estimates is faster than the rate of the oracle learner, it performs asymptotically as if the nuisance functions were known (Kennedy, 2023). However, the convergence rate of the DR-learner will still depend on the convergence rate of its pseudo-outcome regression, which will vary itself depending on the complexity of the CATE.
Despite its desirable properties, the DR-learner can be sensitive to extreme propensity scores, with the IPTWs used in its pseudo-outcomes causing the pseudo-outcomes to grow infinitely large when propensity score estimates are near 0/1. To prevent this, some trim the propensity score estimates, therefore stopping the pseudo-outcomes from growing too large. However, this introduces bias to the propensity score estimates which can propagate through to the CATE estimates themselves. For this reason, we discuss an alternative approach known as infinite-dimensional targeting (iTMLE) (Luedtke et al., 2017; Vansteelandt and Morzywołek, 2023; van der Laan et al., 2024). This technique has been developed for counterfactual outcome prediction (Vansteelandt and Morzywołek, 2023), and tends to have less sensitivity to extreme propensity scores as its targeted learning framework can moderate the impact of outlying propensity score estimates. In this paper we discuss a very similar approach known as the EP-learner (efficient plug-in learner) (van der Laan et al., 2024), which uses iTMLE in the context of causal contrasts, including CATE estimation.
@@
The targeting step then works by regressing the outcomes $Y$ on the sieve basis, $\varphi(X)$, in a weighted linear regression, with an offset $\mu^{A}(Z)$, and weight, $\hat{H}(A,Z)=\frac{A}{\hat{\pi}(Z)}+\frac{1-A}{\left(1-\hat{\pi}(Z)\right)}$, suggested by the EIF of the risk function, eq. (4). After fitting this model, efficient plug-in estimates, $\hat{\mu}^{0*}(Z)$ and $\hat{\mu}^{1*}(Z)$ are obtained by adding/subtracting the estimated linear predictor from this model to the plug-in outcome estimates (See Appendix C for the full algorithm). As the updated outcome predictions are defined such that the sample average of the drift term in eq. (4) converges to 0, the EP-learner achieves the same oracle efficiency properties as the DR-learner. Additionally, by using iTMLE, extreme pseudo-outcomes are less common and the CATE estimates should be more stable (van der Laan et al., 2024).
In addition, both learners require the data which is used in the second stage optimization is i.i.d. To achieve this, sample splitting is used when estimating the nuisance functions, preventing correlations from being introduced into the pseudo-outcomes. In this paper, we implement a K-fold cross-fitting procedure found in Appendix B and C.
@@ 2.3 Estimating the CATE using ML - With missing outcome data
Now we consider how MAR outcome data impacts the existing estimators. To identify the CATE when outcome data is MAR, we require assumptions A1-A3 from Section 2.2 and two additional assumptions on the missingness mechanism: (A4) Outcomes are MAR, $Y\perp\!\!\!\perp C|A,Z$; (A5) Positivity of outcomes being non-missing, $0<P(C=1|A,Z)$, with probability 1. Under assumptions A1-A5, the CATE can be identified as eq. (5) and eq. (6).
@@
The no unmeasured confounding assumption (A2) and MAR assumption (A4) need not be conditional on the same covariates, however, for simplicity, we define both assumptions to be conditional on the same set of covariates, $Z$. Using equations (5) and (6) we see that when all of the covariates in $Z$ are present, the CATE can be estimated using the observed data. Consequently, the existing CATE estimators, such as the T-learner, DR-learner and EP-learner can produce asymptotically unbiased estimates of the CATE by restricting their analyses to complete cases, and adjusting for $Z$ through their outcome models $\mu^{a}(Z)=\operatorname{\mathbb{E}}[Y|A=a,C=1,Z]$, $A\in\{0,1\}$. We refer to this approach as an “available case analysis”, and note that while the outcome regressions are limited to complete cases, the propensity score models should still be estimated using the full dataset.
@@
Available case analyses offer the simplest way of estimating the CATE in the presence of MAR outcome data, however, their asymptotic properties do not assure that they perform well when fit using finite samples. Instead, available case analyses can often be inefficient, as their outcome regressions restrict the population to those with complete cases, with information being ignored for individuals who have a missing outcome. Equally, when the outcome models are fit using ML, they will be prone to over-smoothing in the subsets of the population which experience high levels of outcome missingness. This can cause complex non-linear CATEs to be missed, with these approaches over-smoothing their outcome predictions due to the missing data, or can lead to estimators identifying heterogeneity where non-exists, for instance when the missingness only occurs within one treatment arm. Specific examples of data generating processes (DGPs) where this can occur are presented in Section 4.
@@
Because of these limitations, some authors choose to address missing outcome data by utilizing missing data techniques. A common missing data technique is to impute outcomes, replacing the missing outcomes with outcome predictions gained from an imputation model, $\hat{\operatorname{\mathbb{E}}}[Y|C=1,A,Z]$ (Groenwold et al., 2014; Berrevoets et al., 2023). These are easy to implement when using the existing non-parametric CATE estimators, as the estimators can be run using the imputed, complete dataset. However, as outcome imputations require the estimation of an additional nuisance model, plug-in bias can be introduced, with errors in the outcome imputations propagating through to the estimates of the CATE. Alternatively, other authors suggest addressing missing outcome data by re-weighting observed individuals based on their probability of having a non-missing outcome, $G(A,Z)=P[C=1|A,Z]$ (Robins et al., 1994; Gonzalez Ginestet et al., 2021). These weights can be used to estimate the CATE using a similar approach to the one Kennedy (2023) uses to incorporate IPTWs, weighting observed individuals to create pseudo-outcomes, then regressing the pseudo-outcomes against $X$ to estimate the CATE. However, this estimator, which we will refer to as the IPTW-IPCW estimator, also suffers from plug-in bias, as errors in the IPTWs and IPCWs will propagate through to the CATE estimates. We therefore construct IF based estimators that incorporate these weights naturally, and which offer greater robustness to errors in the missingness predictions $G(A,Z)$.